### v2

@@ 3.3.2 Scan
Table 1: Comparison of different backbones on ImageNet-1K (Deng et al, 2009) classification. indicates that the backbone has undergone additional training using the Token Labeling objective (Wang et al, 2022a). indicates results reproduced by DeiT (Touvron et al, 2021). The best, second-best, and third-best results are bolded and distinguished using red, blue, and green, respectively
Backbone Image Size Params FLOPs Top-1 ACC
CNNs
RegNetY-4G (Radosavovic et al, 2020) $224\times 224$ 20.6 M 4.0 G 80.0
RegNetY-8G (Radosavovic et al, 2020) $224\times 224$ 39.2 M 8.0 G 81.7
RegNetY-16G (Radosavovic et al, 2020) $224\times 224$ 83.6 M 16.0 G 82.9
ConvNeXt-T (Liu et al, 2022b) $224\times 224$ 29.0 M 4.5 G 82.1
ConvNeXt-S (Liu et al, 2022b) $224\times 224$ 50.0 M 8.7 G 83.1
ConvNeXt-B (Liu et al, 2022b) $224\times 224$ 89.0 M 15.4 G 83.8
MambaOut-T (Yu and Wang, 2024) $224\times 224$ 27.0 M 4.5 G 82.7
MambaOut-S (Yu and Wang, 2024) $224\times 224$ 48.0 M 9.0 G 84.1
MambaOut-B (Yu and Wang, 2024) $224\times 224$ 85.0 M 15.8 G 84.2
Transformers
ViT-B/16 (Dosovitskiy et al, 2021) $384\times 384$ 86.0 M 55.4 G 77.9
ViT-L/16 (Dosovitskiy et al, 2021) $384\times 384$ 307.0 M 190.7 G 76.5
DeiT-Ti (Touvron et al, 2021) $224\times 224$ 5.0 M 1.3 G 72.2
DeiT-S (Touvron et al, 2021) $224\times 224$ 22.0 M 4.6 G 79.8
DeiT-B (Touvron et al, 2021) $224\times 224$ 86.0 M 17.5 G 81.8
DeiT-B (Touvron et al, 2021) $384\times 384$ 86.0 M 55.4 G 83.1
Swin-T (Liu et al, 2021) $224\times 224$ 29.0 M 4.5 G 81.3
Swin-S (Liu et al, 2021) $224\times 224$ 50.0 M 8.7 G 83.0
Swin-B (Liu et al, 2021) $224\times 224$ 88.0 M 15.4 G 83.5
Swin-B (Liu et al, 2021) $384\times 384$ 88.0 M 47.0 G 84.5
Wave-ViT-S (Yao et al, 2022) $224\times 224$ 22.7 M 4.7 G 83.9
Wave-ViT-B (Yao et al, 2022) $224\times 224$ 33.5 M 7.2 G 84.8
Wave-ViT-L (Yao et al, 2022) $224\times 224$ 57.5 M 14.8 G 85.5
SVT-H-S (Patro and Agneeswaran, 2023) $224\times 224$ 22.0 M 3.9 G 84.2
SVT-H-B (Patro and Agneeswaran, 2023) $224\times 224$ 32.8 M 6.3 G 85.2
SVT-H-L (Patro and Agneeswaran, 2023) $224\times 224$ 54.0 M 12.7 G 85.7
VOLO-D1 (Yuan et al, 2023) $224\times 224$ 27.0 M 6.8 G 84.2
VOLO-D2 (Yuan et al, 2023) $224\times 224$ 59.0 M 14.1 G 85.2
VOLO-D3 (Yuan et al, 2023) $224\times 224$ 86.0 M 20.6 G 85.4
VOLO-D4 (Yuan et al, 2023) $224\times 224$ 193.0 M 43.8 G 85.7
VOLO-D5 (Yuan et al, 2023) $224\times 224$ 296.0 M 69.0 G 86.1
Mambas
Vim-Ti (Zhu et al, 2024a) $224\times 224$ 7.0 M 1.5 G 76.1
Vim-S (Zhu et al, 2024a) $224\times 224$ 26.0 M 5.1 G 80.5
Mamba®-T (Wang et al, 2024a) $224\times 224$ 9.0 M 5.1 G 77.4
Mamba®-S (Wang et al, 2024a) $224\times 224$ 28.0 M 9.9 G 81.1
Mamba®-B (Wang et al, 2024a) $224\times 224$ 99.0 M 20.3 G 82.9
Mamba®-L (Wang et al, 2024a) $224\times 224$ 341.0 M 55.5 G 83.2
PlainMamba-L1 (Yang et al, 2024a) $224\times 224$ 7.0 M 3.0 G 77.9
PlainMamba-L2 (Yang et al, 2024a) $224\times 224$ 25.0 M 8.1 G 81.6
PlainMamba-L3 (Yang et al, 2024a) $224\times 224$ 50.0 M 14.4 G 82.3
Mamba-2D-S (Li et al, 2024d) $224\times 224$ 24.0 M - 81.7
Mamba-2D-B (Li et al, 2024d) $224\times 224$ 92.0 M - 83.0
VMamba-T (Liu et al, 2024g) $224\times 224$ 31.0 M 4.9 G 82.5
VMamba-S (Liu et al, 2024g) $224\times 224$ 50.0 M 8.7 G 83.6
VMamba-B (Liu et al, 2024g) $224\times 224$ 89.0 M 15.4 G 83.9
FractalMamba-T (Tang et al, 2024b) $224\times 224$ 31.0 M 4.9 G 82.7
Vim-Ti-F(H) (Zhang et al, 2024b) $224\times 224$ 7 M 1.5 G 76.0
Vim-S-F(H) (Zhang et al, 2024b) $224\times 224$ 26 M 5.1 G 80.4
EffVMamba-T (Pei et al, 2024) $224\times 224$ 6.0 M 0.8 G 76.5
EffVMamba-S (Pei et al, 2024) $224\times 224$ 11.0 M 1.3 G 78.7
EffVMamba-B (Pei et al, 2024) $224\times 224$ 33.0 M 4.0 G 81.8
MSVMamba-N (Shi et al, 2024a) $224\times 224$ 7.0 M 0.9 G 77.3
MSVMamba-M (Shi et al, 2024a) $224\times 224$ 12.0 M 1.5 G 79.8
MSVMamba-T (Shi et al, 2024a) $224\times 224$ 33.0 M 4.6 G 82.8
LocalVim-T (Huang et al, 2024e) $224\times 224$ 8.0 M 1.5 G 76.2
LocalVim-S (Huang et al, 2024e) $224\times 224$ 28.0 M 4.8 G 81.2
LocalVMamba-T (Huang et al, 2024e) $224\times 224$ 26.0 M 5.7 G 82.7
LocalVMamba-S (Huang et al, 2024e) $224\times 224$ 50.0 M 11.4 G 83.7
SiMBA-S (Patro and Agneeswaran, 2024b) $224\times 224$ 15.3 M 2.4 G 81.7
SiMBA-B (Patro and Agneeswaran, 2024b) $224\times 224$ 22.8 M 4.2 G 83.5
SiMBA-L (Patro and Agneeswaran, 2024b) $224\times 224$ 36.6 M 7.6 G 84.4
Table 2: Results of object detection and instance segmentation on MS COCO (Lin et al, 2014) $mini$-$val$ using Mask R-CNN (He et al, 2017) 1$\times$ schedule. FLOPs are computed using an input size 1280$\times$800. indicates FLOPs computed using an input size 1333$\times$800. The best, second-best, and third-best results are bolded and distinguished using red, blue, and green, respectively
Backbone Params FLOPs $\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
CNNs
ConvNeXt-T (Liu et al, 2022b) 48.0 M 262.0 G 44.2 66.6 48.3 40.1 63.3 42.8
ConvNeXt-S (Liu et al, 2022b) 70.0 M 348.0 G 45.4 67.9 50.0 41.8 65.2 45.1
MambaOut-T (Yu and Wang, 2024) 43.0 M 262.0 G 45.1 67.3 49.6 41.0 64.1 44.1
MambaOut-S (Yu and Wang, 2024) 65.0 M 354.0 G 47.4 69.1 52.4 42.7 66.1 46.2
MambaOut-B (Yu and Wang, 2024) 100.0 M 495.0 G 47.4 69.3 52.2 43.0 66.4 46.3
Transformers
ViT-Adpt-T (Chen et al, 2023) 28.1 M - 41.1 62.5 44.3 37.5 59.7 39.9
ViT-Adpt-S (Chen et al, 2023) 47.8 M - 44.7 65.8 48.3 39.9 62.5 42.8
ViT-Adpt-B (Chen et al, 2023) 120.2 M - 47.0 68.2 51.4 41.8 65.1 44.9
Swin-T (Liu et al, 2021) 48.0 M 264.0 G 42.2 - - 39.1 - -
Swin-S (Liu et al, 2021) 69.0 M 354.0 G 44.8 - - 40.9 - -
Swin-B (Liu et al, 2021) 107.0 M 496.0 G 46.9 - - 42.3 - -
PVTv2-B2 (Wang et al, 2022b) 45.0 M - 45.3 67.1 49.6 41.2 64.2 44.4
PVTv2-B3 (Wang et al, 2022b) 64.9 M - 47.0 68.1 51.7 42.5 65.7 45.7
PVTv2-B4 (Wang et al, 2022b) 82.2 M - 47.5 68.7 52.0 42.7 66.1 46.1
PVTv2-B5 (Wang et al, 2022b) 101.6 M - 47.4 68.6 51.9 42.5 65.7 46.0
ViL-T (Zhang et al, 2021) 26.9 M 145.6 G 41.4 63.5 45.0 38.1 60.3 40.8
ViL-S (Zhang et al, 2021) 45.0 M 218.3 G 44.9 67.1 49.3 41.0 64.2 44.1
ViL-M (Zhang et al, 2021) 60.1 M 293.8 G 47.6 69.8 52.1 43.0 66.9 46.6
ViL-B (Zhang et al, 2021) 76.1 M 384.4 G 48.6 70.5 53.4 43.6 67.6 47.1
SG-Former-S (Ren et al, 2023) 41.0 M - 47.4 69.0 52.0 42.6 65.9 46.0
SG-Former-M (Ren et al, 2023) 51.0 M - 48.2 70.3 53.1 43.6 66.9 47.0
SG-Former-B (Ren et al, 2023) 95.0 M - 49.2 70.6 54.3 68.2 68.1 47.7
Mambas
PlainMamba-L1 (Yang et al, 2024a) 31.0 M 388.0 G 44.1 64.8 47.9 39.1 61.6 41.9
PlainMamba-L2 (Yang et al, 2024a) 53.0 M 542.0 G 46.0 66.9 50.1 40.6 63.8 43.6
PlainMamba-L3 (Yang et al, 2024a) 79.0 M 696.0 G 46.8 68.0 51.1 41.2 64.7 43.9
VMamba-T (Liu et al, 2024g) 50.0 M 270.0 G 47.4 69.5 52.0 42.7 66.3 46.0
VMamba-S (Liu et al, 2024g) 64.0 M 357.0 G 48.7 70.0 53.4 43.7 67.3 47.0
VMamba-B (Liu et al, 2024g) 108.0 M 485.0 G 49.2 70.9 53.9 43.9 67.7 47.6
FractalMamba-T (Tang et al, 2024b) 50.0 M 270.0 G 47.8 70.0 52.4 42.9 66.6 46.3
EffVMamba-T (Pei et al, 2024) 11.0 M 60.0 G 35.6 57.7 38.0 33.2 54.4 35.1
EffVMamba-S (Pei et al, 2024) 31.0 M 197.0 G 39.3 61.8 42.6 36.7 58.9 39.2
EffVMamba-B (Pei et al, 2024) 53.0 M 252.0 G 43.7 66.2 47.9 40.2 63.3 42.9
MSVMamba-M (Shi et al, 2024a) 32.0 M 201.0 G 43.8 65.8 47.7 39.9 62.9 42.9
MSVMamba-T (Shi et al, 2024a) 53.0 M 252.0 G 46.9 68.8 51.4 42.2 65.6 45.4
LocalVMamba-T (Huang et al, 2024e) 45.0 M 291.0 G 46.7 68.7 50.8 42.2 65.7 45.5
LocalVMamba-S (Huang et al, 2024e) 69.0 M 414.0 G 48.4 69.9 52.7 43.2 66.7 46.5
SiMBA-S (Patro and Agneeswaran, 2024b) 60.0 M 382.0 G 46.9 68.6 51.7 42.6 65.9 45.8
3.3.3 Block
Different combinations of the previously mentioned scanning techniques and the selective SSM transformation form various blocks, which are integral to Mamba-based architectures. In discussing the visual Mamba backbone networks, we provide an overview of these blocks. Fig. 3 illustrates a suite of visual Mamba blocks, including including Vision Mamba (Vim) (Zhu et al, 2024a), Visual State Space (VSS) (Liu et al, 2024g), Efficient Visual State Space (EVSS) (Pei et al, 2024), PlainMamba (Yang et al, 2024a), and LocalMamba (Huang et al, 2024e) blocks. The Mamba (Gu and Dao, 2023) block is also included to facilitate a direct comparison, highlighting the evolutionary design of these blocks in the visual domain. The detailed illustrations of scanning techniques of different blocks and their decoupled results presented in Fig. 3 also validate the logic behind our categorization of scanning techniques. Since the Mamba, Vim, and VSS blocks are extensively utilized in the applications that will be detailed in the subsequent section. For clarity, we simply refer to them as Mamba, Vim and VSS. Their algorithmic details are presented in the Online Resource. Modifications to these blocks are indicated by an asterisk (*).
Table 3: Results of semantic segmentation on ADE20K (Zhou et al, 2019) $val$ using UperNet (Xiao et al, 2018). FLOPs are calculated using an input size 512$\times$2048. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively. MLN: multi-level neck. The best, second-best, and third-best results are bolded and distinguished using red, blue, and green, respectively
Backbone Crop Size Params FLOPs mIoU (SS) mIoU (MS)
CNNs
ConvNeXt-T (Liu et al, 2022b) $512\times 512$ 60.0 M 939.0 G 46.0 46.7
ConvNeXt-S (Liu et al, 2022b) $512\times 512$ 82.0 M 1027.0 G 48.7 49.6
ConvNeXt-B (Liu et al, 2022b) $512\times 512$ 122.0 M 1170.0 G 49.1 49.9
MambaOut-T (Yu and Wang, 2024) $512\times 512$ 54.0 M 938.0 G 47.4 48.6
MambaOut-S (Yu and Wang, 2024) $512\times 512$ 76.0 M 1032.0 G 49.5 50.6
MambaOut-B (Yu and Wang, 2024) $512\times 512$ 112.0 M 1178.0 G 49.6 51.0
Transformers
ViT-Adpt-T (Chen et al, 2023) $512\times 512$ 36.1 M - 42.6 43.6
ViT-Adpt-S (Chen et al, 2023) $512\times 512$ 57.6 M - 46.2 47.1
ViT-Adpt-B (Chen et al, 2023) $512\times 512$ 133.9 M - 48.8 49.7
DeiT-S + MLN (Touvron et al, 2022) $512\times 512$ 58.0 M 1217.0 G 43.8 45.1
DeiT-B + MLN (Touvron et al, 2022) $512\times 512$ 144.0 M 2007.0 G 45.5 47.2
Swin-T (Liu et al, 2021) $512\times 512$ 60.0 M 945.0 G 44.4 45.8
Swin-S (Liu et al, 2021) $512\times 512$ 81.0 M 1039.0 G 47.6 49.5
Swin-B (Liu et al, 2021) $512\times 512$ 121.0 M 1188.0 G 48.1 49.7
SG-Former-S (Ren et al, 2023) $512\times 512$ 52.5 M 989.0 G 49.9 51.5
SG-Former-M (Ren et al, 2023) $512\times 512$ 68.3 M 1114.0 G 51.2 52.1
SG-Former-B (Ren et al, 2023) $512\times 512$ 109.3 M 1304.0 G 52.0 52.7
Mambas
Vim-Ti (Zhu et al, 2024a) $512\times 512$ 13.0 M - 41.0 -
Vim-S (Zhu et al, 2024a) $512\times 512$ 46.0 M - 44.9 -
Mamba®-S (Wang et al, 2024a) $512\times 512$ 56.0 M - 45.3 -
Mamba®-B (Wang et al, 2024a) $512\times 512$ 132.0 M - 47.7 -
Mamba®-L (Wang et al, 2024a) $512\times 512$ 377.0 M - 49.1 -
PlainMamba-L1 (Yang et al, 2024a) $512\times 512$ 35.0 M 174.0 G 44.1 -
PlainMamba-L2 (Yang et al, 2024a) $512\times 512$ 55.0 M 285.0 G 46.8 -
PlainMamba-L3 (Yang et al, 2024a) $512\times 512$ 81.0 M 419.0 G 49.1 -
VMamba-T (Liu et al, 2024g) $512\times 512$ 62.0 M 948.0 G 48.3 48.6
VMamba-S (Liu et al, 2024g) $512\times 512$ 82.0 M 1039.0 G 50.6 51.2
VMamba-B (Liu et al, 2024g) $512\times 512$ 122.0 M 1170.0 G 51.0 51.6
FractalMamba-T (Tang et al, 2024b) $512\times 512$ 62.0 M 948.0 G 48.9 49.8
EffVMamba-T (Pei et al, 2024) $512\times 512$ 14.0 M 230.0 G 38.9 39.3
EffVMamba-S (Pei et al, 2024) $512\times 512$ 29.0 M 505.0 G 41.5 42.1
EffVMamba-B (Pei et al, 2024) $512\times 512$ 65.0 M 930.0 G 46.5 47.3
MSVMamba-M (Shi et al, 2024a) $512\times 512$ 42.0 M 875.0 G 45.1 45.4
MSVMamba-T (Shi et al, 2024a) $512\times 512$ 65.0 M 942.0 G 47.6 48.5
LocalVim-T (Huang et al, 2024e) $512\times 512$ 36.0 M 181.0 G 43.4 44.4
LocalVim-S (Huang et al, 2024e) $512\times 512$ 58.0 M 297.0 G 46.4 47.5
LocalVMamba-T (Huang et al, 2024e) $512\times 512$ 57.0 M 970.0 G 47.9 49.1
LocalVMamba-S (Huang et al, 2024e) $512\times 512$ 81.0 M 1095.0 G 50.0 51.0
SiMBA-S (Patro and Agneeswaran, 2024b) $512\times 512$ 62.0 M 1040.0 G 49.0 49.6
3.4 Experimental Results
In this section, we compare the experimental results of different visual Mamba backbone networks with popular architectures based on CNNs or Transformers on standard computer vision benchmarks.
3.4.1 Image Classification
Image classification performance is compared on ImageNet-1K (Deng et al, 2009) and the Top-1 accuracy results from the respective papers are listed in Table. 1. The results are categorized into three groups based on different architectures: CNNs, Transformers, and Mamba-based models. The best, second-best, and third-best results are bolded and distinguished using red, blue, and green, respectively. To clearly compare different networks, we also plot their performance and computational complexity in Fig. 5.
[…]

### v3

@@ 3.2. Optimization Strategies
Table 3. Comparison of different backbones on ImageNet-1K (Deng et al., 2009) classification at the resolution of $224^{2}$ (Part I). ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. indicates that the backbone has undergone additional training using the Token Labeling objective (Wang et al., 2022a). indicates results are obtained with MESA (Du et al., 2022). indicates results reproduced by DeiT (Touvron et al., 2021). indicates hierarchical structures while others are plain structures. l represents local sampling. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
[…]
Figure 5. Comparative analysis of the performance and computational complexity across various visual backbone architectures, encompassing Convolution-based methods (Radosavovic et al., 2020; Liu et al., 2022; Yu and Wang, 2024), Transformer-based methods (Touvron et al., 2021; Liu et al., 2021; Yao et al., 2022; Patro and Agneeswaran, 2023; Yuan et al., 2023), and Mamba-based methods (Zhu et al., 2024c; Yang et al., 2024a; Liu et al., 2024e; Shi et al., 2024a; Wang et al., 2024j; Xie et al., 2024c; Shi et al., 2024b). The symbol size is proportional to the parameter count of respective model, providing a visual indicator of model scale and complexity. Note that all data are accessed from released academic papers to ensure fairness and credibility.
[…]
3.3. Experimental Results
In this section, we compare the experimental results of different visual Mamba backbone networks with established architectures based on CNNs or Transformers on standard computer vision benchmarks.
3.3.1. Image Classification
Image classification performance is compared on ImageNet-1K (Deng et al., 2009) and the Top-1 accuracy results from the respective papers are listed in Table. 3. To clearly compare different networks, we also plot their performance and computational complexity in Fig. 5. As seen, the performance of all the models improves as model sizes increase, illustrating the trend of enhancing performance through model scaling. It is evident that with similar computational complexity, all visual Mamba networks outperform the CNN-based RegNetY (Radosavovic et al., 2020) and the Transformer-based DeiT (Touvron et al., 2021). Furthermore, many visual Mamba networks surpass the CNN-based MambaOut (Yu and Wang, 2024) and Transformer-based Swin Transformer (Liu et al., 2021). These results demonstrate the superior performance of visual Mamba networks compared to advanced CNNs and common Transformers. However, Transformer-based networks are highly competitive, particularly SVT (Patro and Agneeswaran, 2023), Wave-ViT (Yao et al., 2022), and VOLO (Yuan et al., 2023), which consistently demonstrate superior performance over other networks. Notably, VOLO (Yuan et al., 2023) also exhibits excellent scalability. Currently, despite their efficiency, most Mamba-based networks are limited to small-scale implementations with FLOPs under 21G. While Mamba® (Wang et al., 2024j) and ARM (Ren et al., 2024a) have demonstrated scalability to larger models, there remains a need for performance enhancements. To sum up, visual Mamba networks exhibit promising performance but still underperform advanced Transformer-based networks. Further exploration is also needed to scale visual Mamba networks to larger configurations.
3.3.2. Object Detection
Object detection and instance segmentation performance are compared on the MS COCO (Lin et al., 2014) via Mask R-CNN (He et al., 2017) and the results from the respective papers are listed in Table. 4. The object detection and instance segmentation performance of all the models follows the scaling law as well. It can be observed that with comparable computational complexity, visual Mamba networks surpass the CNN-based ConvNeXt (Liu et al., 2022), except for EfficientVMamba (Pei et al., 2024), which is designed for lightweight purposes. Most visual Mamba networks demonstrate superior performance compared to all CNN-based architectures and a majority of Transformer-based models. Nonetheless, in certain scenarios, their performance is slightly below that of SG-Former (Ren et al., 2023). These results demonstrate that Mamba’s ability to capture long-range dependencies and utilize dynamic weights is advantageous for dense prediction tasks. Currently, Mamba-based models are not specifically tailored for these tasks, indicating potential for future research to explore and enhance their performance.
3.3.3. Semantic Segmentation
Semantic segmentation performance is compared on ADE20K (Zhou et al., 2019) utilizing UperNet (Xiao et al., 2018) and the results from the respective papers are listed in Table. 5. The semantic segmentation performance of different networks is similar to their object detection performance. Many visual Mamba networks outperform all CNN-based networks and most Transformer-based networks. However, their performance lags behind that of SG-Former (Ren et al., 2023). These results further demonstrate Mamba’s efficacy in dense prediction tasks. Further exploration of Mamba-based models specifically designed for these tasks is warranted.
Table 4. Results of object detection and instance segmentation on MS COCO (Lin et al., 2014) $mini$-$val$ using Mask R-CNN (He et al., 2017) 1$\times$ schedule. ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. FLOPs are computed using an input size 1280$\times$800. indicates FLOPs computed using an input size 1333$\times$800. indicates hierarchical structures while others are plain structures. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
Backbone Type Params FLOPs $\mathbf{AP^{b}}$ $\mathbf{AP^{b}_{50}}$ $\mathbf{AP^{b}_{75}}$ $\mathbf{AP^{m}}$ $\mathbf{AP^{m}_{50}}$ $\mathbf{AP^{m}_{75}}$
ConvNeXt-T (Liu et al., 2022) Conv 48.0 M 262.0 G 44.2 66.6 48.3 40.1 63.3 42.8
MambaOut-T (Yu and Wang, 2024) Conv 43.0 M 262.0 G 45.1 67.3 49.6 41.0 64.1 44.1
Swin-T (Liu et al., 2021) Attn 48.0 M 264.0 G 42.2 - - 39.1 - -
PVTv2-B2 (Wang et al., 2022b) Attn 45.0 M - 45.3 67.1 49.6 41.2 64.2 44.4
ViL-S (Zhang et al., 2021) Attn 45.0 M 218.3 G 44.9 67.1 49.3 41.0 64.2 44.1
SG-Former-S (Ren et al., 2023) Attn 41.0 M - 47.4 69.0 52.0 42.6 65.9 46.0
[…]
VMamba-T (Liu et al., 2024e) Mamba 50.0 M 270.0 G 47.4 69.5 52.0 42.7 66.3 46.0
FractalMamba-T (Tang et al., 2024c) Mamba 50.0 M 270.0 G 47.8 70.0 52.4 42.9 66.6 46.3
EffVMamba-S (Pei et al., 2024) Mamba 31.0 M 197.0 G 39.3 61.8 42.6 36.7 58.9 39.2
V2M-S (Wang et al., 2024m) Mamba 50 M - 47.6 69.4 52.2 42.9 66.5 46.3
GlobalMamba-T (Wang et al., 2024p) Mamba 50 M - 47.6 69.4 52.5 42.9 66.5 46.0
QuadMamba-T (Xie et al., 2024c) Mamba 30 M 213 G 42.3 64.6 46.2 38.8 61.6 41.4
MSVMamba-M (Shi et al., 2024a) Mamba 32.0 M 201.0 G 43.8 65.8 47.7 39.9 62.9 42.9
LocalVMamba-T (Huang et al., 2024d) Mamba 45.0 M 291.0 G 46.7 68.7 50.8 42.2 65.7 45.5
GroupMamba (Shaker et al., 2024) Mamba 40 M 279 G 47.6 69.8 52.1 42.9 66.5 46.3
Spatial-Mamba-T (Xiao et al., 2024b) Mamba 46 M 261 G 47.6 69.6 52.3 42.9 66.5 46.2
VSSD-M (Shi et al., 2024b) Mamba 33 M 220 G 45.4 67.5 49.8 41.3 64.5 44.6
VSSD-T (Shi et al., 2024b) Mamba 44 M 265 G 46.9 69.4 51.4 42.6 66.4 45.9
ConvNeXt-S (Liu et al., 2022) Conv 70.0 M 348.0 G 45.4 67.9 50.0 41.8 65.2 45.1
MambaOut-S (Yu and Wang, 2024) Conv 65.0 M 354.0 G 47.4 69.1 52.4 42.7 66.1 46.2
ViT-Adpt-S (Chen et al., 2023) Attn 47.8 M - 44.7 65.8 48.3 39.9 62.5 42.8
Swin-S (Liu et al., 2021) Attn 69.0 M 354.0 G 44.8 - - 40.9 - -
PVTv2-B3 (Wang et al., 2022b) Attn 64.9 M - 47.0 68.1 51.7 42.5 65.7 45.7
ViL-M (Zhang et al., 2021) Attn 60.1 M 293.8 G 47.6 69.8 52.1 43.0 66.9 46.6
SG-Former-M (Ren et al., 2023) Attn 51.0 M - 48.2 70.3 53.1 43.6 66.9 47.0
[…]
VMamba-S (Liu et al., 2024e) Mamba 64.0 M 357.0 G 48.7 70.0 53.4 43.7 67.3 47.0
EffVMamba-B (Pei et al., 2024) Mamba 53.0 M 252.0 G 43.7 66.2 47.9 40.2 63.3 42.9
V2M-B (Wang et al., 2024m) Mamba 70 M - 48.9 70.2 53.6 43.8 67.5 47.1
GlobalMamba-S (Wang et al., 2024p) Mamba 70 M - 49.0 70.5 53.5 43.9 67.5 47.0
QuadMamba-S (Xie et al., 2024c) Mamba 55 M 301 G 46.7 69.0 51.3 42.4 65.9 45.6
MSVMamba-T (Shi et al., 2024a) Mamba 53.0 M 252.0 G 46.9 68.8 51.4 42.2 65.6 45.4
LocalVMamba-S (Huang et al., 2024d) Mamba 69.0 M 414.0 G 48.4 69.9 52.7 43.2 66.7 46.5
SiMBA-S (Patro and Agneeswaran, 2024b) Mamba 60.0 M 382.0 G 46.9 68.6 51.7 42.6 65.9 45.8
Spatial-Mamba-S (Xiao et al., 2024b) Mamba 63 M 315 G 49.2 70.8 54.2 44.0 67.9 47.5
VSSD-S (Shi et al., 2024b) Mamba 59 M 325 G 48.4 70.1 53.1 43.5 67.2 47.1
MambaOut-B (Yu and Wang, 2024) Conv 100.0 M 495.0 G 47.4 69.3 52.2 43.0 66.4 46.3
ViT-Adpt-B (Chen et al., 2023) Attn 120.2 M - 47.0 68.2 51.4 41.8 65.1 44.9
PVTv2-B4 (Wang et al., 2022b) Attn 82.2 M - 47.5 68.7 52.0 42.7 66.1 46.1
PVTv2-B5 (Wang et al., 2022b) Attn 101.6 M - 47.4 68.6 51.9 42.5 65.7 46.0
Swin-B (Liu et al., 2021) Attn 107.0 M 496.0 G 46.9 - - 42.3 - -
ViL-B (Zhang et al., 2021) Attn 76.1 M 384.4 G 48.6 70.5 53.4 43.6 67.6 47.1
SG-Former-B (Ren et al., 2023) Attn 95.0 M - 49.2 70.6 54.3 68.2 68.1 47.7
[…]
VMamba-B (Liu et al., 2024e) Mamba 108.0 M 485.0 G 49.2 70.9 53.9 43.9 67.7 47.6
GlobalMamba-B (Wang et al., 2024p) Mamba 108 M - 49.3 71.4 54.2 44.2 68.4 47.7
Spatial-Mamba-B (Xiao et al., 2024b) Mamba 115 M 494 G 50.4 71.8 55.3 45.1 69.1 49.1
Table 5. Results of semantic segmentation on ADE20K (Zhou et al., 2019) $val$ at the resolution of $512^{2}$ using UperNet (Xiao et al., 2018). ‘Conv’, ‘Attn’, and ‘Mamba’ denote architectures based on CNN, Transformer, and Mamba, respectively. FLOPs are calculated using an input size 512$\times$2048. ‘SS’ and ‘MS’ denote single-scale and multi-scale testing, respectively. MLN: multi-level neck. indicates hierarchical structures while others are plain structures. The best results are marked with bold and underline, the second-best with bold only, and the third-best with underline only.
Backbone Type Params FLOPs mIoU (SS) mIoU (MS)
ViT-Adpt-T (Chen et al., 2023) Attn 36.1 M - 42.6 43.6
Vim-Ti (Zhu et al., 2024c) Mamba 13.0 M - 41.0 -
PlainMamba-L1 (Yang et al., 2024a) Mamba 35.0 M 174.0 G 44.1 -
EffVMamba-T (Pei et al., 2024) Mamba 14.0 M 230.0 G 38.9 39.3
QuadMamba-T (Xie et al., 2024c) Mamba 40 M 886 G 44.3 45.1
VSSD-M (Shi et al., 2024b) Mamba 42 M 893 G 45.6 46.0
ConvNeXt-T (Liu et al., 2022) Conv 60.0 M 939.0 G 46.0 46.7
MambaOut-T (Yu and Wang, 2024) Conv 54.0 M 938.0 G 47.4 48.6
ViT-Adpt-S (Chen et al., 2023) Attn 57.6 M - 46.2 47.1
DeiT-S + MLN (Touvron et al., 2022) Attn 58.0 M 1217.0 G 43.8 45.1
Swin-T (Liu et al., 2021) Attn 60.0 M 945.0 G 44.4 45.8
SG-Former-S (Ren et al., 2023) Attn 52.5 M 989.0 G 49.9 51.5
[…]
Mamba®-S (Wang et al., 2024j) Mamba 56.0 M - 45.3 -
PlainMamba-L2 (Yang et al., 2024a) Mamba 55.0 M 285.0 G 46.8 -
VMamba-T (Liu et al., 2024e) Mamba 62.0 M 948.0 G 48.3 48.6
FractalMamba-T (Tang et al., 2024c) Mamba 62.0 M 948.0 G 48.9 49.8
EffVMamba-S (Pei et al., 2024) Mamba 29.0 M 505.0 G 41.5 42.1
V2M-S (Wang et al., 2024m) Mamba 62 M - 48.2 49.0
GlobalMamba-T (Wang et al., 2024p) Mamba 62 M - 48.1 49.0
QuadMamba-S (Xie et al., 2024c) Mamba 62 M 961 G 47.2 48.1
MSVMamba-M (Shi et al., 2024a) Mamba 42.0 M 875.0 G 45.1 45.4
LocalVim-T (Huang et al., 2024d) Mamba 36.0 M 181.0 G 43.4 44.4
LocalVMamba-T (Huang et al., 2024d) Mamba 57.0 M 970.0 G 47.9 49.1
MambaVision-T (Hatamizadeh and Kautz, 2024) Mamba 55 M 945 G 46.6 -
GroupMamba-T (Shaker et al., 2024) Mamba 49 M 955 G 48.6 49.2
Spatial-Mamba-T (Xiao et al., 2024b) Mamba 57 M 936 G 48.6 49.4
VSSD-T (Shi et al., 2024b) Mamba 53 M 941 G 47.9 48.7
Backbone Type Params FLOPs mIoU (SS) mIoU (MS)
ConvNeXt-S (Liu et al., 2022) Conv 82.0 M 1027.0 G 48.7 49.6
MambaOut-S (Yu and Wang, 2024) Conv 76.0 M 1032.0 G 49.5 50.6
Swin-S (Liu et al., 2021) Attn 81.0 M 1039.0 G 47.6 49.5
SG-Former-M (Ren et al., 2023) Attn 68.3 M 1114.0 G 51.2 52.1
[…]
VMamba-S (Liu et al., 2024e) Mamba 82.0 M 1039.0 G 50.6 51.2
EffVMamba-B (Pei et al., 2024) Mamba 65.0 M 930.0 G 46.5 47.3
V2M-B (Wang et al., 2024m) Mamba 82 M - 50.8 51.3
[…]